OpenStudy (anonymous):

convert the radian measure to degree measure. Use the value of π found on a calculator and round answers to two decimals places. -π/3

OpenStudy (anonymous):

there are 360 degrees = 2*pi radians or 180 = pi radians This is now just a conversion from radians to degrees

OpenStudy (anonymous):

\[-\frac{\pi }{3}*\frac{180}{\pi }=-60 \text{ degrees}\]

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

@juantweaver or @robtobey can you help with a similar problem. 5